# MultiBLiMP 1.0: A Massively Multilingual Benchmark of Linguistic Minimal Pairs

### v1

@@ 4.3 Inflection
To form the ungrammatical counterpart of a sentence extracted from UD, we inflect a specific word to a different feature that makes the sentence ungrammatical. For example, to create a minimal pair for subject-verb number agreement for The boy walks, we inflect walks for number=Sing to create The boy walk. While we focus here on gender, person, and number, this can work for any agreement feature annotated on the verb.
Our inflection procedure has two stages. We start by defining the feature to inflect for, with a candidate word from the filtered treebank of § 4.1. The first stage is to find the corresponding lemma and morphological features of this word in the UD and UM databases, based on its form and the features that were annotated in the UD tree. This may yield multiple candidate lemmas and features, depending on the level of detail of the UD features. In the second stage, we then find all matching rows containing opposite values of the feature that we inflect for. The inflected forms of these matching rows are the inflection candidates that we use to create the minimal pairs. By considering all opposite values we may create multiple pairs from a single sentence, e.g. He is [..] → He are/am [..].
@@ 4.4 Dataset Balancing
To level data imbalances, we limit the number of minimal pairs per language condition to 100 items. We obtain these 100 items using a weighted downsampling procedure that also balances sentence-level features. The features we incorporate in this procedure are: subject form, verb form, subject-verb distance, all attractors congruent, and all attractors incongruent. The latter two indicate whether the intervening material between subject and verb contained items that had a congruent or incongruent feature with respect to the inflected feature. We define the sampling probability of a minimal pair as the inverse joint probability of its features, assuming feature independence:
$P(x)=\prod_{i}P(x_{i})\hskip 28.45274ptQ(x)=\frac{P(x)^{-1}}{\sum_{i}P(\mathbf{x}_{i})^{-1}}$
The feature probability $P(x_{i})$ is expressed as the relative frequency of the feature. By sampling from $Q$ (without replacements), we can balance all features simultaneously.
@@
We run our minimal pair creation pipeline for two agreement phenomena and three features: subject-verb and subject-participle agreement for number, person, and gender. Within each condition we create sub-conditions based on the inflected feature (e.g. SG $\rightarrow$ PL, 2 $\rightarrow$ 3, etc.), and the order of subject and verb. Our procedure results in 128,321 minimal pairs across 101  languages. The total number of minimal pairs we obtain before balancing is 1.4 million. We provide an aggregated overview of the number of languages per condition in Table 2, and a sample of pairs in § A.2. Subject-verb number agreement is the most common agreement type, with 90 languages, whereas we find subject-participle person agreement for only 3 languages. We also provide a detailed breakdown of the language families that are currently covered in MultiBLiMP 1.0 in Figure 3. There is a strong Indo-European (IE) bias present in our minimal pairs, caused both by the over-representation of such languages in UD, as well as our choice to focus on agreement phenomena, which tend to be more present in IE languages. In future iterations of MultiBLiMP, we intend to broaden our coverage by focusing on a more diverse set of phenomena.
Subject-Verb Subject-Participle Resources Languages
Model Size Variant # P G # P G Low Mid High GF Aya EU Eng All #best
@@
8B it 82.6 86.6 86.8 88.4 94.8 86.5 75.7 88.9 95.6 88.3 94.6 91.4 99.4 85.6 0
8B tülu3 81.7 86.2 85.5 87.3 94.9 85.6 74.7 88.1 95.3 87.6 94.0 90.9 98.6 84.9 0
70B base 87.4 91.2 91.1 92.1 97.5 91.2 81.1 93.9 97.8 92.6 97.1 95.5 99.0 90.2 7
Aya 8B it 76.9 82.4 81.8 80.8 93.3 79.3 68.3 82.3 96.3 83.3 95.7 88.0 99.0 80.4 0
32B it 82.9 87.8 88.1 87.4 95.1 87.0 75.7 89.4 97.7 89.0 97.3 92.8 98.4 86.4 1
Gemma3 4B base 82.8 87.2 86.0 88.8 93.3 86.0 71.1 93.0 96.4 89.2 95.3 94.6 98.2 85.8 0
4B it 73.4 79.1 76.9 79.3 87.9 76.9 63.9 82.7 88.2 80.0 86.6 85.1 93.4 77.1 0
12B base 85.4 90.2 89.6 91.1 95.5 90.3 75.8 95.4 97.7 92.0 96.8 96.6 99.0 88.8 0
12B it 79.2 84.6 83.1 85.1 93.0 83.8 69.5 89.4 92.6 85.9 91.4 90.9 95.6 82.9 0
27B base 87.2 91.1 91.3 92.8 96.6 91.7 78.3 96.3 98.0 93.2 97.4 97.1 98.6 90.2 3
27B it 82.7 87.0 86.3 88.7 95.8 86.7 73.1 92.3 94.4 88.9 93.7 93.3 96.0 85.8 0
@@
GPT2 1.6B xl 61.1 67.8 55.9 63.2 53.2 57.4 62.8 61.2 68.4 63.5 66.0 64.0 98.3 63.4 0
Goldfish 500M base 92.4 95.3 92.2 95.2 98.2 90.9 88.0 95.6 95.9 93.8 95.2 95.8 96.4 93.8 14
Table 3: Average accuracies per LM, split out for different phenomena and language subsets. GF is the language subset of Goldfish languages; Aya the subset of Aya languages; EU the subset of EuroLLM languages; Eng the performance on English. The best performing model per category is denoted in boldface, but we exclude Goldfish models from this ranking as they are only evaluated on a subset of languages. #best denotes the number of languages for which this model was significantly better than the others.
6 LM Evaluation Setup
6.1 Language Models
We evaluate the following LMs, accessed via HuggingFace (Wolf et al., 2020): Llama3 Grattafiori et al. (2024) in sizes 8B and 70B; base and post-trained. We also evaluate Tülu3 Lambert et al. (2025), which is a post-trained model on the Llama3 base models. Aya-expanse Dang et al. (2024b), in sizes 8B and 32B, has been post-trained on a set of 23 languages, 18 of which are in MultiBLiMP 1.0. We include Gemma3 Team (2025) in sizes 4B, 12B and 27B; base and post-trained, which were pre-trained on a balanced multilingual distribution of 140 languages. OLMo2 OLMo et al. (2025) in size 32B; base and post-trained. EuroLLM Martins et al. (2024), for size 9B, which is pre-trained on 35 (mostly) European languages, 32 of which are present in MultiBLiMP 1.0.
Monolingual LMs
The aforementioned LMs are all multilingual, i.e. trained on a mixture of different languages. To place their performance into perspective, we also evaluate the Goldfish suite (Chang et al., 2024), a collection of monolingual LMs each trained on the same amount of data. This allows us to control for language frequency differences, which is not possible for the pre-trained LMs. In our experiments we consider the models trained on 1GB of data, or the full models for languages with less than 1GB available data. All models have 500M parameters. Goldfish models exist for 70 out of the 101 languages in MultiBLiMP. As a monolingual English baseline, we also evaluate on GPT2-xl (Radford et al., 2019).
6.2 Training Corpus Language Distribution
In our experiments we connect model performance to the language distribution of the training data. Since the training corpora of most LLMs (except OLMo2) are not publicly available, we estimate this distribution based on the language frequencies of Kargaran et al. (2024), which were computed on a 3.9T token split of the Common Crawl corpus. Common Crawl provides a good reflection of the language distribution of the web-scraped data that is at the core of many LLM training corpora.
@@ 6.3 Metrics
Following prior work (Marvin and Linzen, 2018) we measure the grammaticality of the predictions of the LMs based on the probability it assigns to a grammatical sentence versus an ungrammatical one. We chose to evaluate on the sentence-level, and not at the position of the key items, as has been done previously (Linzen et al., 2016; Jumelet and Hupkes, 2018), as this yields the same metric for both subject-verb orders. We use two metrics in our experiments. The accuracy score is based on the number of items for which the LM assign a higher probability to the grammatical sentence ($s^{+}$):
@@
Figure 4: Average accuracy score per language, plotted for Llama3-8B against various other models.
Figure 5: The $\beta$ coefficient estimates of OLS models fitted on $\Delta$ scores of Llama3-70B and Gemma3-27B.
@@ 7.1 General Performance
The average accuracies of all models are presented in Table 5. Results are split into various categories: the six agreement phenomena; three language splits based on resourcedness; the language subsets of the Goldfish, Aya, and EuroLLM models; performance on English and the overall average. The low-, mid-, and high-resource splits are based on the Common Crawl language frequencies: low-resource for the least frequent 60% languages, mid-resource for the languages in the 60–90% frequency range, and high-resource for the most frequent 10% languages. We also denote the number of languages for which a model was significantly better than all other models. [^3] We conduct a McNemar test between the best and second-best model based on the item-level binary grammaticality judgments, with $p<0.1$ as significance threshold.
Llama3-70B-base and Gemma3-27B-base perform best overall, both scoring 90.2% on average. These two models also score best on all the categories, with the Llama model scoring better on low-resource languages, and Gemma better on mid- and high-resource languages. Gemma3-27B-base outperforms the Aya and EuroLLM models on their specific language subsets, validating the claim of Team (2025) that its better balance of languages leads to strong multilingual performance. Interestingly, both these models are still being outperformed by the 500M Goldfish model series. The Goldfish models are significantly better on 14 languages, whereas Llama3-70B and Gemma3-27B are best for 7 and 3 languages, respectively. We present the best model per language in § B.2.
Performance is strongly driven by model size: for all model types, performance increases with number of parameters. Furthermore, instruction tuning has a detrimental effect on performance, which is especially prevalent for the Gemma3 models. For all models, performance increases with language frequency, demonstrating much of the variance in linguistic ability can be explained by data disparity. The performance of GPT2-xl shows that LLMs must be endowed with strong multilingual abilities to perform well on MultiBLiMP. We show each model’s performance per language in Appendix B.3.
@@ 7.2 Cross-Model Comparisons
We plot the language-level performance of various models against each other in Figure 4. Each model is compared against Llama3-8B to investigate different factors driving performance.
In Figure 4a we consider the impact of model size, comparing the 8B model to the 70B model. The average performance went up from 86.9% to 90.2% for this model, driven by gradual improvements: no language improved by more than 10%, but almost all languages improved to some degree. We also conduct a McNemar test between the two models, with $p<0.05$, resulting in Llama3-70B being significantly better for 48 languages, and none the other way around.
Next, we consider the impact of language-specific post- and pre-training in Fig. 4b and 4c, comparing Llama3-8B to Aya-8B and EuroLLM-9B. We have highlighted the specific languages on which Aya has been post-trained, and EuroLLM has been pre-trained. Languages outside this selection do not appear to benefit from cross-lingual transfer, as both Aya and EuroLLM perform worse than Llama3-8B in the languages they were not trained on, and McNemar testing shows Llama3-8B is significantly better than Aya for 49 languages, and better than EuroLLM for 31 languages. Language-specific pre-training has a far stronger effect on linguistic ability than post-training: EuroLLM significantly outperforms Llama3-8B on 19 of its 32 languages, whereas Aya only outperforms it on 4 out of its 18 languages.
Finally, we look at the impact of monolingual vs multilingual language modelling in Figure 4d. Goldfish models significantly outperform Llama3-8B for 39 out of 70 languages, whereas Llama only does so for 7 languages. This warrants a more detailed investigation in future work, to assess if this difference is solely driven by differences in data frequency and data quality, or by the multilingual LLM having to handle all these languages at once.
@@ 7.3 What Factors Drive Performance?
We set up a linear modelling experiment in which we predict the $\Delta$ scores (§6.3) of a few models based on various factors that we hypothesize to be driving model judgments. We include the following fixed effects in our regression model: the six agreement phenomena as categorical variables; subject-verb distance; perplexity of the grammatical sentence; sentence length; subword delta expressing the difference in the number of subwords for the grammatical and ungrammatical key item; two binary features denoting if the intervening material between subject and verb contained congruent or incongruent attractors; and the language frequency of §6.2. To interpret this model we plot the $\beta$ coefficients for the standardized factors, fitted for the two best-performing LLMs.
As shown in Fig. 5, the phenomenon coefficients show that person agreement yields high $\Delta$ scores (i.e. more confidence in detecting the grammatical sentence). This can be due to the fact that first and second person subjects tend to be closed-class pronouns, whereas gender and number agreement requires acquisition of open-class subject features. Surprisingly, subject-verb distance does not have a negative impact on $\Delta$: one might expect agreement to be more challenging for long-distance dependencies. An increase in sentence perplexity strongly decreases $\Delta$: the more surprised a model is by a sentence, the less likely it is to make the right grammaticality judgment. Another important factor with a negative effect is the subword delta: if the correct form of the verb is split into more subwords than the incorrect form, the model is more likely to make a wrong judgment. As we expected, incongruent attractors result in a decrease in $\Delta$, while congruent attractors boost it. Finally, language frequency has a positive effect on performance. A more detailed breakdown of this effect is shown in Fig. 1, where we highlight the languages that over- and underperform with respect to a fitted logistic curve. An interesting question for future work would be to investigate the extent to which the variance in performance is driven by typological features.

### v2

@@ 4.3 Inflection
To form the ungrammatical counterpart of a sentence extracted from UD, we inflect a specific word to a different feature that makes the sentence ungrammatical. For example, to create a minimal pair for subject-verb number agreement for The boy walks, we re-inflect the verb walk with number=Plur to create The boy walk. While we focus here on gender, person, and number, this can work for any agreement feature annotated on the verb.
Our inflection procedure has two stages. We start by defining the feature to inflect for, with a candidate word from the filtered treebank of § 4.1. The first stage is to find the corresponding lemma and morphological features of this word in the UD and UM databases, based on its form and the features that were annotated in the UD tree. This may yield multiple candidate lemmas and features, depending on the level of detail of the UD features. In the second stage, we then find all matching rows containing contrasting values of the feature that we inflect for. The inflected forms of these matching rows are the inflection candidates that we use to create the minimal pairs. By considering all opposite values we may create multiple pairs from a single sentence, e.g. He is [..] → He are/am [..].
@@ 4.4 Dataset Balancing
To level data imbalances, we limit the number of minimal pairs per agreement condition to 100 items. We obtain these 100 items using a weighted downsampling procedure that also balances sentence-level features. The features we incorporate in this procedure are: subject form, verb form, subject-verb distance, all attractors congruent, and all attractors incongruent. The last two features indicate whether the intervening material between subject and verb contained items that had a congruent or incongruent feature with respect to the agreement feature that we re-inflect (e.g. The keys to the cabinet are). We define the sampling probability $Q$ of a minimal pair as the inverse of the joint probability of its features $P(x)$, assuming feature independence:
$P(x)=\prod_{i}P(x_{i})\hskip 28.45274ptQ(x)=\frac{P(x)^{-1}}{\sum_{i}P({x}_{i})^{-1}}$
The feature probability $P(x_{i})$ is expressed as the relative frequency of the feature. By sampling from $Q$ (without replacements), we can balance all features simultaneously. This results in a dataset that is as balanced as possible for multiple features, ensuring both lexical diversity and an even spread of easier and harder sentences.
@@
We run our minimal pair creation pipeline for two agreement phenomena and three features: subject-verb and subject-participle agreement for number, person, and gender. Within each condition we create sub-conditions based on the inflected feature (e.g. SG $\rightarrow$ PL, 2 $\rightarrow$ 3, etc.), and the order of subject and verb. Our procedure results in 128,321 minimal pairs across 101  languages. The total number of minimal pairs we obtain before balancing is 1.4 million. We provide an aggregated overview of the number of languages per condition in Table 2, and a sample of pairs in § A.2. Subject-verb number agreement is the most common agreement type, with 90 languages, whereas we find subject-participle person agreement for only 3 languages. We present the number of minimal pairs per language in Figure 4. We also provide a detailed breakdown of the language families that are currently covered in MultiBLiMP 1.0 in Figure 3. There is a strong Indo-European (IE) bias present in our minimal pairs, caused both by the over-representation of such languages in UD, as well as our choice to focus on agreement phenomena, which tend to be more present in IE languages. In future iterations of MultiBLiMP, we intend to broaden our coverage by focusing on a more diverse set of phenomena.
Subject-Verb Subject-Participle Resources Language Subset
Model Size Version N P G N P G Low Mid High GF Aya EU Eng All #best
@@
70B base 87.4 91.2 91.1 92.1 97.5 91.2 81.1 93.9 97.8 92.6 97.1 95.5 99.0 90.2 2
70B chat 86.5 90.3 90.3 91.7 97.2 90.6 80.3 93.1 96.9 91.9 96.2 94.6 98.3 89.3 0
Aya 32B chat 82.9 87.8 88.1 87.4 95.1 87.0 75.7 89.4 97.7 89.0 97.3 92.8 98.4 86.4 1
Gemma3 27B base 87.2 91.1 91.3 92.8 96.6 91.7 78.3 96.3 98.0 93.2 97.4 97.1 98.6 90.2 3
27B chat 82.7 87.0 86.3 88.7 95.8 86.7 73.1 92.3 94.4 88.9 93.7 93.3 96.0 85.8 0
@@
Goldfish 125M 92.4 95.3 92.2 95.2 98.2 90.9 88.0 95.6 95.9 93.8 95.2 95.8 96.4 93.8 14
Table 3: Average accuracies per LLM, split out for different phenomena and language subsets. N stands for Number, P for Person, G for Gender. GF is the language subset of Goldfish languages; Aya the subset of Aya languages; EU the subset of EuroLLM languages; Eng the performance on English. The best performing model per category is denoted in boldface, but we exclude Goldfish models from this ranking as they are only evaluated on a subset of languages. #best denotes the number of languages for which this model was significantly better than the others.
6 LLM Evaluation Setup
@@ 6.1 Metrics
Following prior work (Marvin and Linzen, 2018) we measure the grammaticality of the predictions of a model based on the probability it assigns to a grammatical sentence versus an ungrammatical one.666 The best way to evaluate instruction-tuned LLMs on minimal pairs and grammaticality remains a topic of debate. ‘Meta-linguistic’ evaluation assesses grammaticality judgments not based on direct probability comparisons, but instead prompts the model to pick the grammatical sentence. While Dentella et al. (2023) found GPT3’s abilities lacking using this approach, Hu et al. (2024) show that direct evaluation may yield better results, and Song et al. (2025) show that meta-linguistic evaluation depends strongly on model size, and outperforms direct evaluation for larger models. We choose to evaluate on the sentence-level, and not at the position of the key items, as has been done previously (Linzen et al., 2016; Jumelet and Hupkes, 2018), as this yields the same metric for both subject-verb word orders. We use two metrics in our experiments. The accuracy score is based on the number of items for which the model assigns a higher probability to the grammatical sentence ($s^{+}$):
@@
6.2 Language Models
We evaluate the following 42 LLMs, accessed through HuggingFace (Wolf et al., 2020). For all models, we evaluate both their base version (the model after pre-training), and chat version (the model after post-training).
  • Llama3 Grattafiori et al. (2024) in sizes 1B, 3B, 8B and 70B. We also evaluate Tülu3 Lambert et al. (2025), which is post-trained on the Llama3 8B base model.
  • Aya-expanse Dang et al. (2024), in sizes 8B and 32B, which have been post-trained on a set of 23 languages, 18 of which are in MultiBLiMP 1.0.
  • Gemma3 Kamath et al. (2025) in sizes 4B, 12B and 27B, which were pre-trained on a balanced multilingual distribution of 140 languages.
  • Qwen3 (Yang et al., 2025) in sizes 0.6B, 1.7B, 4B, 8B, and 14B.
  • OLMo2 OLMo et al. (2025) in sizes 1B, 7B, 14B, and 32B.
  • EuroLLM Martins et al. (2024), for sizes 1.7B and 9B, which are pre-trained on 35 (mostly) European languages, 32 of which are present in MultiBLiMP 1.0.
Monolingual Models
The aforementioned LLMs are all multilingual, i.e. trained on a mixture of different languages. To place their performance into perspective, we also evaluate the Goldfish suite (Chang et al., 2024b), a collection of monolingual models each trained on the same amount of data. This allows us to control for language frequency differences, which is not possible for the pre-trained LLMs. In our experiments we consider the models trained on 1GB of data, or the full models for languages with less than 1GB available data. All models have 500M parameters. Goldfish models exist for 70 out of the 101 languages in MultiBLiMP.
6.3 Training Corpus Language Distribution
In our experiments we connect model performance to the language distribution of the training data. Since the training corpora of most LLMs are not publicly available, we estimate this distribution based on the language frequencies of Kargaran et al. (2024), which were computed on a 3.9T token split of the Common Crawl corpus. Common Crawl provides a good reflection of the language distribution of the web-scraped data that is at the core of many LLM training corpora.
Figure 5: The impact of model size (in number of parameters on a log scale) against overall MultiBLiMP accuracy for the Llama3, Qwen3, Gemma3, and OLMo2 model families.
Figure 6: Average accuracy score per language, plotted for Llama3-8B against various other models. The Aya and EuroLLM languages are a subset of the Certain Agreement class, and are highlighted for visualization purposes.
@@ 7.1 General Performance
We report the average accuracy of the largest model for each model family in Table 5. We consider results separately for subject-verb and subject-participle agreement, and then split further by feature between Number, Person, and Gender. We also report results for language subgroups split based on the Common Crawl language frequencies: low-resource for the least frequent 60% languages, mid-resource for the languages in the 60–90% frequency range, and high-resource for the most frequent 10% languages. Finally, we report results for the language subsets of the Goldfish, Aya, and EuroLLM models; performance on English and the overall average. We also denote the number of languages for which a model was significantly better than all other models. [^7] We conduct a McNemar test between the best and second-best model based on the item-level binary grammaticality judgments, with $p<0.1$ as significance threshold.
Llama3-70B-base and Gemma3-27B-base perform best overall, both scoring 90.2% on average. These two models also score best on most of the categories, with the Llama model scoring better on low-resource languages, and Gemma better on mid- and high-resource languages. For all models, performance increases with language frequency, demonstrating much of the variance in linguistic ability can be explained by data disparity. Gemma3-27B-base outperforms the Aya and EuroLLM models on their specific language subsets, validating the claim of Kamath et al. (2025) that its better balance of languages leads to strong multilingual performance. Interestingly, both these models are still being outperformed by the Goldfish model series. The Goldfish models are significantly better on 14 out of 101 languages, whereas Llama3-70B and Gemma3-27B are best for 2 and 3 languages, respectively. We present the best model per language in § B.2, and a breakdown of per-language performance in Appendix B.3.
7.2 Impact of Model Properties
We investigate how model properties drive performance, focusing on model size and post-training. For this analysis we make use of the full set of all 42 LLMs described in §6.2. We fitted a linear regression model to predict the average accuracy of the LLMs, using model size as a continuous variable, model family as a categorical fixed effect, and a binary indicator for post-training status (post-trained vs. base).
The fitted model yields an $R^{2}$ of 0.936. Model size has a significantly positive impact on performance ($\beta=4.30$; $t=15.88$). Post-training, however, has significantly negative impact on performance ($\beta=-3.29$; $t=-6.47$). This finding is in line with Chirkova and Nikoulina (2024), who report that instruction tuning (a stage of post-training) can have a negative effect on multilingual fluency in LLMs. [^8] We experimented with different chat templates for the chat version of the Gemma3 models, but did not observe any improvement. We leave more elaborate ‘prompt engineering’ open for future work. We visualize this trend in Figure 5 for the four largest model families: performance increases with size, and post-trained models consistently underperform their base version.
@@ 7.3 Cross-Model Comparisons
While the previous experiment demonstrates that model size and post-training have an impact on performance, it remains unclear how individual languages are impacted by this. To investigate this, we plot the language-level performance of various models against each other in Figure 6.
We consider the impact of language-specific post- and pre-training in Fig. 6a) and 6b), comparing the chat version of Llama3-8B to Aya-8B and the base version of Llama3-8B to EuroLLM-9B. We have highlighted the specific languages on which Aya has been post-trained, and EuroLLM has been pre-trained. Languages outside this selection do not appear to benefit from cross-lingual transfer, as both Aya and EuroLLM perform worse than Llama3-8B in the languages they were not trained on, and McNemar testing shows Llama3-8B is significantly better than Aya for 43 languages, and better than EuroLLM for 31 languages. This suggests that language-specific pre-training has a stronger effect on linguistic ability than post-training: EuroLLM significantly outperforms Llama3-8B on 19 of its 32 languages (59%), whereas Aya only outperforms it on 8 out of its 18 languages (44%). Further research in a setting with better control over the training data is needed to identify more precisely the data properties that drive multilingual grammar acquisition.
Next, we consider the impact of model size in Figure 6c), by comparing the 8B model to the 70B model. The average performance goes up from 86.9% to 90.2% for this model, driven by gradual improvements: no language improves by more than 10%,, but almost all languages improve to some degree. We also conduct a McNemar test between the two models, with $p<0.05$, resulting in Llama3-70B being significantly better for 48 languages, and none the other way around.
Finally, we look at the impact of monolingual vs multilingual language modelling in Figure 6d. Goldfish models significantly outperform Llama3-8B for 39 out of 70 languages, whereas Llama only does so for 7 languages. This warrants a more detailed investigation in future work, to assess if this difference is solely driven by differences in data frequency and data quality, or by the multilingual LLM having to handle all these languages in a shared representation space (Chang et al., 2024a).
@@ 7.4 What Factors Drive Performance?
We set up a linear modelling experiment in which we predict the $\Delta$ scores (§6.1) of a few models based on various factors that we hypothesize to be driving model judgments. We include the following fixed effects in our regression model: the six agreement types as categorical variables; subject-verb distance; perplexity of the grammatical sentence; sentence length; subword delta expressing the difference in the number of subwords for the grammatical and ungrammatical key item; two binary features denoting if the intervening material between subject and verb contained congruent or incongruent attractors; and the language frequency of §6.3. To interpret this model, we plot the $\beta$ coefficients for the standardized factors, fitted for the two best-performing LLMs (Llama3-70B and Gemma3-27B), and the Goldfish models.
Results
As shown in Fig. 7, the agreement type coefficients show that person agreement yields high $\Delta$ scores (i.e. more confidence in detecting the grammatical sentence). This can be due to the fact that first and second person subjects tend to be closed-class pronouns, whereas gender and number agreement requires acquisition of open-class subject features. Surprisingly, subject-verb distance does not have a negative impact on $\Delta$: one might expect agreement to be more challenging for long-distance dependencies. An increase in sentence perplexity strongly decreases $\Delta$: the more surprised a model is by a sentence, the less likely it is to make the right grammaticality judgment.
Figure 7: The $\beta$ coefficient estimates of OLS models fitted on $\Delta$ scores of Llama3-70B, Gemma3-27B, and Goldfish.
Another important factor with a negative effect is the subword delta: if the correct form of the verb is split into more subwords than the incorrect form, the model is more likely to make a wrong judgment. As we expected, incongruent attractors result in a decrease in $\Delta$, while congruent attractors boost it. Finally, language frequency has a positive effect on the performance of the LLMs, whereas for Goldfish (which controls for frequency) this is not the case. This is the only factor for which the Goldfish coefficient is opposite to the LLMs, for all other factors it is similar. A more detailed breakdown of impact of language frequency on Gemma3-27B is shown in Fig. 1, where we highlight the languages that over- and underperform with respect to a fitted logistic curve. An interesting question for future work would be to investigate the extent to which the variance in performance is driven by typological features.